What mass of ice -140c Will Be needed to cool 200cm2 Of an orange drink (essential water) from 250C to 1000c

1
Expert's answer
2021-07-12T12:12:56-0400
"mc_i(T_0-T_i)+mc_w(T-T_0)+mL\\=Mc_w(T_w-T)\\
m(2090)(14)+m(4200)(25)\\+(334000)m=(0.2)(4200)(75)\\m=0.135\ kg"